Is there an instance when a linear equation is not a function? ...
can you give me an example....create equation of a nonlinear function with tweo inputs

Answer by stanbon(57250)

Is there an instance when a linear equation is not a function? ...
can you give me an example
x=1 is linear but is not a function.
-------------------------------------
....create equation of a nonlinear function with tweo inputs
Let (1,2) and (3,4) be two points on a line. Find the equation.
slope = 4-2/3-1 = 2/2 = 1
--------------------------
Form: y = mx+b
2 = 1*1+b
b = 1
EQUATION:
y = x + 1
===============
Cheers,
Stan H.